OpenStudy (anonymous):

Solve for x. x^2 + 6x = 0

OpenStudy (anonymous):

@skullpatrol gooooooo awayyy like seriously? stalk someone else's questions...

OpenStudy (anonymous):

This question is done exactly the same way your previous questions were solved. What seems to be the problem? The method is exactly the same.

OpenStudy (anonymous):

cool.

OpenStudy (skullpatrol):

I'm sorry if you think I'm "stalking" you :(

OpenStudy (skullpatrol):

Are you ready to learn @xcaseynoel ?

OpenStudy (usukidoll):

1. factor out the x.

OpenStudy (skullpatrol):

Yes, 1. factor out the x is the first step Do you know why it works @xcaseynoel ?

OpenStudy (usukidoll):

-____-....meh... then solve for each x. yay the end.

OpenStudy (usukidoll):

is this just a regular problem or one of those complete the square problems? that's what I'm wondering.

OpenStudy (skullpatrol):

Because of the distributive property of multiplication with respect to addition: $$ab+ac=a(b+c)$$ $$x^2+6x= x(x+6)$$

OpenStudy (anonymous):

\[(-6\pm6)/2\] hence x=0 or x=-6

OpenStudy (skullpatrol):

@xcaseynoel Does that make sense?

OpenStudy (usukidoll):

ok just the regular problem..if that's the case then solve for each x which is 0 and -6

OpenStudy (anonymous):

x= (−6±6)/2 hence x=0 or x=-6

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

@skullpatrol see i used the formula for quadratic equation by which i was showing a better way to solve the question quickly. The factoring method here may take longer time and by this formula one can calculate such questions mentally.

OpenStudy (skullpatrol):

@leena1996 but factoring is a valuable skill to learn how to do fast, no?

OpenStudy (anonymous):

@skullpatrol Factoring is a good technique to learn but each question requires a different approach and some questions are not possible with factoring so one needs to apply other methods which can fetch them the answer without wasting time and learninig new techniques.